OpenStudy (anonymous):

Please Help!!!!! Patty's Pet Company charges $0.55 per pound to ship pet supplies. Part A: Write an equation to determine the total cost, c, to ship p pounds of pet supplies. Use your equation to determine the cost of shipping 2 pounds of pet supplies. (6 points) Part B: If the company reduces the cost to ship pet supplies by 0.05 per pound, write an equation to determine the total cost, c, to ship p pounds of pet supplies with the reduced cost. (4 points)

OpenStudy (anonymous):

part A: 0.55p=c (0.55)(2)=$1.1 part B:0.05p=c (0.05)(2)=$0.01

OpenStudy (anonymous):

Part A) Equation (c=0.55p) (c=0.55*2) c=1.1 Part B) Equation (c=0.50p) (c=0.50*2) c=1.1

OpenStudy (anonymous):

my bad part b (0.50)(2)=$1.00

OpenStudy (anonymous):

Mine too 1.00

OpenStudy (anonymous):

my bad part b (0.50)(2)=$1.00

OpenStudy (anonymous):

Do you want to know how we came with the equation and answer

OpenStudy (anonymous):

yes please